# Learning an Optimal Assortment Policy under Observational Data

### v1

@@ 4.2 Theoretical Results
For any integer $K$ satisfying $N\geq 5K$, there exists a set of observed $K$-sized assortments $D_{S}=\{S_{k}\}_{k=1}^{n}\subset[N]^{n}$, a reward vector $\bm{r}\in[0,1]^{N}$ and a class of attraction parameters $\mathcal{V}\subset[0,1]^{N}$, so that for any algorithm $\mathcal{A}$ that takes as input $D_{S}$ and
@@
For any $K$ and $N\geq 5K$, when the reward vector $\bm{r}=\bm{1}_{N}$, there exists a set of observed $K$-sized assortments $D_{S}=\{S_{k}\}_{k=1}^{n}\subset[N]^{n}$ and a set of problem parameters $\mathcal{V}\subset[0,1]^{N}$. So that for any algorithm $\mathcal{A}$ that takes as input $D_{S}$ and
@@ 5.1 Construction of Hard Problem Instances
    Optimal items $\mathcal{N}_{\text{opt}}$: A $K$-sized subset of $[4K]$ such that each item in this subset has attraction value of $1/K+\epsilon$ and revenue of $1$, with $\epsilon=\frac{1}{25\sqrt{n_{\min}}}.$
@@
To define $\mathcal{V},$ we introduce the following packing number result under the distance $\Delta(S,S^{\prime}):=\lvert(S\cup S^{\prime})\setminus(S\cap S^{\prime})\rvert$ defined over $\mathcal{S}_{K}:$
@@
For sufficiently large $K$, there exists a collection $\mathcal{F}$ of $K$-sized subsets of $[4K]$ so that $\Delta(S,S^{\prime})\geq K/2,\forall S,S^{\prime}\in\mathcal{F}$ and $\log\lvert\mathcal{F}\rvert\geq K/20.$
With the above $K/2$-packing collection $\mathcal{F}$ of $[4K]$, we define $\mathcal{V}$ as the set consisting of all $(\bm{v},\bm{r})$ pairs induced by all possible partitions described above with $\mathcal{N}_{\text{opt}}\in\mathcal{F}$. Note that the partition is fully determined by the choice of $\mathcal{N}_{\text{opt}}$, and therefore, the size of $\mathcal{V}$ satifies $\log\lvert\mathcal{V}\rvert=\log\lvert\mathcal{F}\rvert\geq K/20$. Since there is a one-to-one correspondence between $(\bm{r},\bm{v})\in\mathcal{V}$ and the partition, we use the notation $\mathcal{N}_{\text{opt}}(\bm{r},\bm{v})$, $\mathcal{N}_{c}(\bm{r},\bm{v})$, and $\mathcal{N}_{0}(\bm{r},\bm{v})$ for $(\bm{r},\bm{v})\in\mathcal{V}$ to denote the optimal, competitive, and significantly sub-optimal item sets, respectively, under the parameters $\bm{r},\bm{v}$.
@@ 5.2 Applying Fano’s Lemma
Firstly, by our construction , we have the following lower bound result regarding the sub-optimality gap via cardinality-based set distance $\Delta(\cdot,\cdot),$ defined as
$\Delta(S,S^{\prime}):=\lvert(S\cup S^{\prime})\setminus(S\cap S^{\prime})\rvert.$
@@
we have by $\log M=\log\lvert\mathcal{V}\rvert\geq K/20,$ Lemma 5.3 implies that
$\min_{\pi}\max_{(\bm{r},\bm{v})\in\mathcal{V}}{\mathbb{E}}[\Delta(\mathcal{N}_{\text{opt}}(\bm{r},\bm{v}),\pi(D))]\gtrsim K\left(\frac{1}{25}-\frac{\sum_{(\bm{r},\bm{v})\in\mathcal{V}}\sum_{(\bm{r}^{\prime},\bm{v}^{\prime})\in\mathcal{V}}D(\mathbb{P}_{\bm{r},\bm{v}}\lVert\mathbb{P}_{\bm{r}^{\prime},\bm{v}^{\prime}})}{K\lvert\mathcal{V}\rvert^{2}}\right).$
@@ 5.3 Upper Bounding the KL-Divergence
Lemma 5.4 indicates that with the selection $\epsilon=\frac{1}{25\sqrt{n}_{\min}},$ we have $\max_{(\bm{r},\bm{v}),(\bm{r}^{\prime},\bm{v}^{\prime})}D(\mathbb{P}_{\bm{r},\bm{v}}\lVert\mathbb{P}_{\bm{r}^{\prime},\bm{v}^{\prime}})\leq K/50,$ which then leads to
$K\left(\frac{1}{25}-\frac{\sum_{(\bm{r},\bm{v})\in\mathcal{V}}\sum_{(\bm{r}^{\prime},\bm{v}^{\prime})\in\mathcal{V}}D(\mathbb{P}_{\bm{r},\bm{v}}\lVert\mathbb{P}_{\bm{r}^{\prime},\bm{v}^{\prime}})}{K\lvert\mathcal{V}\rvert^{2}}\right)\geq\frac{K}{50},$
@@ Appendix B Proof of Auxiliary Results in Section 5
Without loss of generality, we suppose $K$ is a even integer, then notice that for every $S\in\mathcal{S}_{K},$ its $K/2$-neighborhood under $\Delta,$ defined as
$\delta(S,K/2):=\{S^{\prime}\subset[4K]:\lvert S^{\prime}\rvert=K,\Delta(S,S^{\prime})\leq K/2\},$
satisfies $\lvert\delta(S,K/2)\rvert\leq 1+\sum_{m=1}^{K/2}\binom{K}{m}\binom{3K}{K-m}.$ If we consider the maximal $K/2$-packing $\mathcal{F}$ of $\mathcal{S}_{K}$ under $\Delta$, then we should have
$\left(1+\sum_{m=1}^{K/2}\binom{K}{m}\binom{3K}{K-m}\right)\lvert\mathcal{F}\rvert\geq\binom{4K}{K}.$
Since otherwise, there will exists a $K$-sized set that does not lies in all $K/2$ neighborhoods of $S\in\mathcal{F},$ a contradiction to $\mathcal{F}$ is maximal.
Now for any $1\leq m\leq K/2,$ we have
$\frac{\binom{4K}{K}}{\binom{K}{m}\binom{3K}{K-m}}\geq\frac{\binom{4K}{K}}{\binom{K}{K/2}\binom{3K}{K/2}}=\frac{(4K)!}{K!(3K)!}\bigg{/}\left(\frac{K!(3K)!}{((K/2)!)^{3}(5K/2)!}\right)=(4K)!(K/2)!(5K/2)!\left(\frac{(K/2)!}{(3K)!K!}\right)^{2},$
using Stirling’s approximation, where $n!\sim\sqrt{2\pi n}(n/e)^{n},$ we have then for sufficiently large $K$, there exists some absolute constant $C>0$ so that the above lower bound turns to
$(4K)!(K/2)!(5K/2)!\left(\frac{(K/2)!}{(3K)!K!}\right)^{2}$ $\geq CK^{1/2}\cdot\frac{(4K)^{4K}(K/2)^{3K/2}(5K/2)^{5K/2}}{(3K)^{6K}K^{2K}}$
$\geq CK^{1/2}\left(\frac{4^{4}\cdot(5/2)^{5/2}\cdot(1/2)^{3/2}}{3^{6}}\right)^{K}\geq CK^{1/2}(6/5)^{K},$
that then leads to $\log\lvert\mathcal{F}\rvert\geq\frac{1}{2}K\log(6/5)\geq K/20$ for sufficiently large $K$. ∎
@@ C.1 Modifications on the hard-instance construction
    For items in $\mathcal{N}_{\text{opt}}$, we modify the attraction value to $\frac{1}{K}+\epsilon^{\prime}$ with $\epsilon^{\prime}=\frac{1}{\sqrt{25Kn_{\min}}}.$
@@
$\min_{\pi}\max_{(\bm{r},\bm{v})\in\mathcal{V}}{\mathbb{E}}[\Delta(\mathcal{N}_{\text{opt}}(\bm{r},\bm{v}),\pi(D))]\geq K\left(\frac{1}{25}-\frac{\sum_{(\bm{r},\bm{v})\in\mathcal{V}}\sum_{(\bm{r}^{\prime},\bm{v}^{\prime})\in\mathcal{V}}D(\mathbb{P}_{\bm{r},\bm{v}}\lVert\mathbb{P}_{\bm{r}^{\prime},\bm{v}^{\prime}})}{K\lvert\mathcal{V}\rvert^{2}}\right).$
@@ C.2 Upper bounds of KL-divergence
Now with (C.2) and $\epsilon=\frac{1}{25\sqrt{Kn_{\min}}},$ we have $\max_{(\bm{r},\bm{v}),(\bm{r}^{\prime},\bm{v}^{\prime})}D(\mathbb{P}_{\bm{r},\bm{v}}\lVert\mathbb{P}_{\bm{r}^{\prime},\bm{v}^{\prime}})\leq K/50,$ which then leads to
$K\cdot\left(\frac{1}{25}-\frac{\sum_{(\bm{r},\bm{v})\in\mathcal{V}}\sum_{(\bm{r}^{\prime},\bm{v}^{\prime})\in\mathcal{V}}D(\mathbb{P}_{\bm{r},\bm{v}}\lVert\mathbb{P}_{\bm{r}^{\prime},\bm{v}^{\prime}})}{K\lvert\mathcal{V}\rvert^{2}}\right)\geq\frac{K}{50},$

### v2

@@ 4.2 Theoretical Results
For any sufficiently large integers $N,K$ satisfying $N\geq 5K$, there exists a set of observed $K$-sized assortments $D_{S}=\{S_{k}\}_{k=1}^{n}\subset[N]^{n}$, a reward vector $\bm{r}\in[0,1]^{N}$ and a class of attraction parameters $\mathcal{V}\subset[0,1]^{N}$, so that for any algorithm $\mathcal{A}$ that takes as input $D_{S}$ and
@@
For any sufficiently large integers $N,K$ such that $N\geq 5K$, when the reward vector $\bm{r}=\bm{1}_{N}$, there exists a set of observed $K$-sized assortments $D_{S}=\{S_{k}\}_{k=1}^{n}\subset[N]^{n}$ and a set of problem parameters $\mathcal{V}\subset[0,1]^{N}$. So that for any algorithm $\mathcal{A}$ that takes as input $D_{S}$ and
@@ 5.1 Construction of Hard Problem Instances
    Optimal items $\mathcal{N}_{\text{opt}}$: A $K$-sized subset of $[4K]$ such that each item in this subset has attraction value of $1/K+\epsilon$ and revenue of $1$, with $\epsilon<\frac{1}{4K}$ to be determined later.
@@
To define $\mathcal{V},$ we introduce the following packing number result under the distance
$\Delta(S,S^{\prime}):=\lvert(S\cup S^{\prime})\setminus(S\cap S^{\prime})\rvert$ (5.1)
defined over $\mathcal{S}_{K}:$
@@
There exists some absolute constants $C_{0},C_{1}>0$ so that when $K>C_{0}$, there exists a collection $\mathcal{F}$ of $K$-sized subsets of $[4K]$ so that $\Delta(S,S^{\prime})\geq\lfloor K/4\rfloor,\forall S,S^{\prime}\in\mathcal{F}$ and $\log\lvert\mathcal{F}\rvert\geq C_{1}K.$
With the above $K/4$-packing collection $\mathcal{F}$ of $[4K]$, we define $\mathcal{V}$ as the set consisting of all $(\bm{v},\bm{r})$ pairs induced by all possible partitions described above with $\mathcal{N}_{\text{opt}}\in\mathcal{F}$. Note that the partition is fully determined by the choice of $\mathcal{N}_{\text{opt}}$, and therefore, the size of $\mathcal{V}$ satisfies $\log\lvert\mathcal{V}\rvert=\log\lvert\mathcal{F}\rvert\geq C_{1}K$. Since there is a one-to-one correspondence between $(\bm{r},\bm{v})\in\mathcal{V}$ and the partition, we use the notation $\mathcal{N}_{\text{opt}}(\bm{r},\bm{v})$, $\mathcal{N}_{c}(\bm{r},\bm{v})$, and $\mathcal{N}_{0}(\bm{r},\bm{v})$ for $(\bm{r},\bm{v})\in\mathcal{V}$ to denote the optimal, competitive, and significantly sub-optimal item sets, respectively, under the parameters $\bm{r},\bm{v}$.
@@ 5.2 Applying Fano’s Lemma
Firstly, by our construction , we have the following lower bound result regarding the sub-optimality gap via cardinality-based set distance $\Delta(\cdot,\cdot),$ defined in (5.1)
@@
we have by $\log M=\log\lvert\mathcal{V}\rvert\geq C_{1}K,$ Lemma 5.3 implies that when $K\geq C_{2}:=\max\{C_{0},C_{1}^{-1}2\log 2\}$
$\min_{\pi}\max_{(\bm{r},\bm{v})\in\mathcal{V}}{\mathbb{E}}[\Delta(\mathcal{N}_{\text{opt}}(\bm{r},\bm{v}),\pi(D))]\gtrsim K\left(\frac{1}{2}-\frac{\sum_{(\bm{r},\bm{v})\in\mathcal{V}}\sum_{(\bm{r}^{\prime},\bm{v}^{\prime})\in\mathcal{V}}D(\mathbb{P}_{\bm{r},\bm{v}}\lVert\mathbb{P}_{\bm{r}^{\prime},\bm{v}^{\prime}})}{C_{1}K\lvert\mathcal{V}\rvert^{2}}\right).$
@@ 5.3 Upper Bounding the KL-Divergence
Lemma 5.4 indicates that with the selection $\epsilon=\sqrt{\frac{C_{1}}{20n_{\min}}},$ we have $\max_{(\bm{r},\bm{v}),(\bm{r}^{\prime},\bm{v}^{\prime})}D(\mathbb{P}_{\bm{r},\bm{v}}\lVert\mathbb{P}_{\bm{r}^{\prime},\bm{v}^{\prime}})\leq C_{1}K/4,$ which then leads to
$K\left(\frac{1}{2}-\frac{\sum_{(\bm{r},\bm{v})\in\mathcal{V}}\sum_{(\bm{r}^{\prime},\bm{v}^{\prime})\in\mathcal{V}}D(\mathbb{P}_{\bm{r},\bm{v}}\lVert\mathbb{P}_{\bm{r}^{\prime},\bm{v}^{\prime}})}{C_{1}K\lvert\mathcal{V}\rvert^{2}}\right)\geq\frac{K}{4},$
@@ Appendix B Proof of Auxiliary Results in Section 5
We assume without loss of generality that $K/8$ is an integer, otherwise we can replace $K/4,K/8$ by $\lfloor K/4\rfloor,\lfloor K/8\rfloor$ all the analysis still hold. Now for every $K$-sized subset $S$ of $[4K]$, define its $K/4$-neighborhood under $\Delta$ as
$\delta(S,K/4):=\{S^{\prime}\subset[4K]:\lvert S^{\prime}\rvert=K,\Delta(S,S^{\prime})\leq K/4\}.$
We have then denote $\mathcal{F}$ as the maximal $K/2$-packing set consists of $K$-sized subsets of $[4K]$, we have
$\lvert\delta(S,K/4)\rvert\cdot\lvert\mathcal{F}\rvert\geq\binom{4K}{K}.$
Since otherwise there will exists some $K$-sized set does not lies in any $K/4$ neighborhood of $S\in\mathcal{F},$ a contradiction to $\mathcal{F}$ is maximal.
Now it suffice to provide lower bound for $\binom{4K}{K}/\lvert\delta(S,K/4)\rvert:$
Upper bound for $\delta(S,K/2)$.
For any $S,S^{\prime}\subset[4K]$ with $\lvert S\rvert=\lvert S^{\prime}\rvert=K$, $\Delta(S,S^{\prime})$ must be even and for $m\leq K/4$ we have
$\lvert S\cap S^{\prime}\rvert=K-m\iff\lvert S\cup S^{\prime}\rvert=K+m\iff\Delta(S,S^{\prime})=2m,$
as a result, we have
$\lvert\delta(S,K/4)\rvert$ $\leq 1+\sum_{m=1}^{K/8}\lvert\{S^{\prime}:\lvert S^{\prime}\rvert=K,\lvert S\cap S^{\prime}\rvert=m\}\rvert$
$=1+\sum_{m=1}^{K/8}\binom{K}{m}\binom{3K}{K-m}$
$\leq\frac{K}{2}\cdot\max_{1\leq m\leq K/8}\underbrace{\binom{K}{m}\binom{3K}{K-m}}_{:=f(m)}.$
Now since for every $1\leq m\leq K/8,$ we have
$\frac{f(m+1)}{f(m)}=\frac{\binom{K}{m+1}\binom{3K}{K-(m+1)}}{\binom{K}{m}\binom{3K}{K-m}}=\frac{(K-m)^{2}}{(m+1)(2K+m+1)},$
thus for every $m<K/4,$ we have
$\frac{(K-m)^{2}}{(m+1)(2K+m+1)}>\frac{{9}K^{2}/{16}}{\frac{K}{4}\cdot\frac{9}{4}K}=1,$
i.e. the maximal of $f(\cdot)$ on $\{1,\dots,K/8\}$ is attained at $m=K/8.$
As a result, we get
$\frac{K}{2}\cdot\lvert\mathcal{F}\rvert$ $\geq\frac{\binom{4K}{K}}{f(\frac{K}{8})}=\frac{\binom{4K}{K}}{\binom{K}{\frac{K}{8}}\binom{3K}{\frac{7K}{8}}}=\frac{(4K)!}{K!(3K)!}\bigg{/}\bigg{(}\frac{K!(3K)!}{(\frac{7K}{8})!(\frac{K}{8})!\cdot(\frac{7K}{8})!(\frac{17K}{8})!}\bigg{)}.$
Now applying Stirling’s approximation, where there exists some absolute constant $c>0$ so that
$e^{-c/n}\leq\frac{n!}{\sqrt{2\pi n}(n/e)^{n}}\leq e^{c/n},\quad\forall n\geq 1,$ (B.1)
we have there exists some absolute constants $c_{1},c_{2},c^{\prime}>0$ so that
$\frac{(4K)!}{K!(3K)!}\bigg{/}\bigg{(}\frac{K!(3K)!}{(\frac{7K}{8})!(\frac{K}{8})!\cdot(\frac{7K}{8})!(\frac{17K}{8})!}\bigg{)}$
$=(4K)!(\frac{K}{8})!(\frac{17K}{8})!\cdot\bigg{(}\frac{(7K/8)!}{K!(3K)!}\bigg{)}^{2}$
$\geq_{\text{(i)}}c_{1}e^{-c_{2}/K}\sqrt{K}\frac{(4K)^{4K}(K/8)^{K/8}(17K/8)^{17K/8}(7K/8)^{7K/4}}{(3K)^{6K}K^{2K}}$
$\geq c^{\prime}\sqrt{K}\bigg{(}\underbrace{\frac{4^{4}(1/8)^{1/8}(17/8)^{17/8}(7/8)^{7/4}}{3^{6}}}_{\geq\frac{26}{25}}\bigg{)}^{K}.$
Where (i) is by (B.1). Thus we can get
$\lvert\mathcal{F}\rvert\geq\frac{2c^{\prime}}{\sqrt{K}}\cdot(\frac{26}{25})^{K},$
as a result, there exists some absolute number $C_{0}>0$ so that when $K>C_{0},$ we have then
$\frac{2C^{\prime}}{\sqrt{K}}(\frac{26}{25})^{K/2}>1\implies\lvert\mathcal{F}\rvert\geq(\frac{26}{25})^{K/2}\implies\log\lvert\mathcal{F}\rvert$ $\geq\underbrace{\frac{\log(\frac{26}{25})}{2}}_{:=C_{1}}K,$
as desired. ∎
@@ C.1 Modifications on the hard-instance construction
    For items in $\mathcal{N}_{\text{opt}}$, we modify the attraction value to $\frac{1}{K}+\epsilon^{\prime}$ with $\epsilon^{\prime}<\frac{1}{4K}$ to be specified later.
@@
$\min_{\pi}\max_{(\bm{r},\bm{v})\in\mathcal{V}}{\mathbb{E}}[\Delta(\mathcal{N}_{\text{opt}}(\bm{r},\bm{v}),\pi(D))]\geq K\left(\frac{1}{2}-\frac{\sum_{(\bm{r},\bm{v})\in\mathcal{V}}\sum_{(\bm{r}^{\prime},\bm{v}^{\prime})\in\mathcal{V}}D(\mathbb{P}_{\bm{r},\bm{v}}\lVert\mathbb{P}_{\bm{r}^{\prime},\bm{v}^{\prime}})}{C_{1}K\lvert\mathcal{V}\rvert^{2}}\right).$
@@ C.2 Upper bounds of KL-divergence
Now with (C.2) and set $\epsilon^{\prime}=\sqrt{\frac{C_{1}}{20Kn_{\min}}},$ we have $\max_{(\bm{r},\bm{v}),(\bm{r}^{\prime},\bm{v}^{\prime})}D(\mathbb{P}_{\bm{r},\bm{v}}\lVert\mathbb{P}_{\bm{r}^{\prime},\bm{v}^{\prime}})\leq C_{1}K/4,$ which then leads to
$K\cdot\left(\frac{1}{2}-\frac{\sum_{(\bm{r},\bm{v})\in\mathcal{V}}\sum_{(\bm{r}^{\prime},\bm{v}^{\prime})\in\mathcal{V}}D(\mathbb{P}_{\bm{r},\bm{v}}\lVert\mathbb{P}_{\bm{r}^{\prime},\bm{v}^{\prime}})}{C_{1}K\lvert\mathcal{V}\rvert^{2}}\right)\geq\frac{K}{4},$